OpenStudy (anonymous):

find an nth-degree polynomial function with real coefficients satisfying the give conditions n=3;-4 and 6+5i are zero; f(1)=250

OpenStudy (anonymous):

So it's a degree 3.

OpenStudy (anonymous):

If \(6+5i\) is a root than so is \(6-5i\)

OpenStudy (anonymous):

So you know it is of the form\[ c(x-(-4))(x-(6+5i))(x-(6-5i)) = c(x+4)(x^2-12x+61) \]

OpenStudy (anonymous):

You can solve for \(c\) by using: \[ 250=c((1)+4)((1)^2-12(1)+61) \]

OpenStudy (anonymous):

hey how can u solve (x-(6+5i))(x-(6-5i))=(x^2-12x+61) can u explain more thank

OpenStudy (tkhunny):

Solve by multiplication and comparison. The coefficients on the two resulting polynomials must be equal.